OpenStudy (anonymous):

If f(x) = 2x2 - 10, find f(5).

OpenStudy (anonymous):

f(x) = 2x^2 - 10 f(5) = 2*5^2-10 = 2 * 25 - 10 = 40

OpenStudy (anonymous):

??